What would be the intercepts of the line that contains pair of points:
" (-1,-4) and (6,10) " ?

Answer by TimothyLamb(4379)

points on line:
(-1, -4)
(6, 10)
---
use two-point form:
y - y1 = ( (y2 - y1)/(x2 - x1) )(x - x1)
---
y + 4 = ( (10 + 4)/(6 + 1) )(x + 1)
y + 4 = ( (14)/(7) )(x + 1)
y + 4 = 2(x + 1)
y + 4 = 2x + 2
y = 2x - 2
---
intercepts:
---
a is the x-intercept:
y = 2x - 2
0 = 2x - 2
2 = 2x
x = 1
a = (1, 0)
---
b is the y-intercept:
y = 2x - 2
y = 2(0) - 2
y = -2
b = (0, -2)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php